If Sally is 3 years more than 4 times her youngest child. Together they ad up to 48. How old is Sally ? How old is her youngest daughter?

Answer by josh_jordan(263)

To solve, we need to change this word problem into an equation. First, it says that Sally is 3 years more than 4 times her youngest child. Let's use the letter S to represent Sally and y to represent her youngest child. If Sally is 3 years more than 4 times her youngest child, we can set this up as:

S = 3 + 4y

Next, we are told that the sum of their ages is 48. So,

S + y = 48

We can replace S with 3 + 4y, giving us

3 + 4y + y = 48

Next, combine 4y + y, giving us

3 + 5y = 48

Next, we need to isolate the y variable. To do this, we can subtract 3 from both sides of the equation:

3 - 3 + 5y = 48 - 3, giving us

5y = 45

Next, we can divide both sides of the equation by 5:

5y/5 = 45/5, which gives us

y = 9

Therefore, her youngest child is 9 years old. To find Sally's age, we can replace the y in S = 3 + 4y with 9:

S = 3 + 4(9), which gives us

S = 39

Therefore, Sally is 39 years old.

Final Answer: Sally is 39 years old and her youngest daughter is 9 years old.